OpenStudy (anonymous):

Hey I have my final exam today and I was looking over my review and found a problem I would like explained if anyone could assist with that. Solve the equation: x^2 + 7x - 60 = 0 (The answers are x = -12 and x=5) Just a breakdown of how to work it out would be great!

OpenStudy (anonymous):

well by the looks of it you just do -12^2 + 7x5 or times 5 - 60 = 0 (tho i dont know what ^ means im only in 6th)

OpenStudy (anonymous):

\[(x-5)(x+12)=x^2+7x-60=0\] multiply x times x to get x squared x times 12 to get 12x multiply -5 times x=-5x and -5 times 12=-60 put these together x^2+7x-60

OpenStudy (anonymous):

Thanks Jamaica :)

OpenStudy (anonymous):

@RyanM.Freeman ^ means to the power of. It's just easier to use than to go and find a squared symbol.